Given the values of x and y, find an equation of variation in which y varies directly as x. y=8, when x=5.
Answer by tutorcecilia(2152)

y=kx [formula for direct variation]
(8)=k(5) [plug-in the variables and solve for the k-term]
8=5k
8/5=5k/5
8/5=k
.
y=(8/5)x [this is the equation of variation]